OpenStudy (anonymous):

Sound travels at a rate of 330 meters per second. If a crash occurs 7,920 meters from where you are standing, how many seconds does it take for you to hear the sound?

OpenStudy (anonymous):

WELCOME TO \[\huge OPENSTUDY \] If u have any wonders and doubt on \[\huge OPENSTUDY \] consult @ParthKohli @lgbasallote @mathslover @rebeccaskell94 @ash2326 Okay have a best time here Regards...

OpenStudy (cwrw238):

=7920 / 330 seconds

OpenStudy (anonymous):

@Ron.mystery I don't know you and I appreciate you trying to help but I think @cshalvey wanted us to avoid this method of welcoming new users as it is ovewhelming and distracts from their question. also @cwrw238 you really shouldn't just give answers we are here to help them learn, not supply them with just the answers :)

OpenStudy (anonymous):

speed = distance / time time = distance / speed time = 7920/330 time = 24 seconds

OpenStudy (anonymous):

please there was no \[\huge\ AMBASSADOR\] here so I took a role there as i \[\huge LOVVVEEEE \] this PLACE A LOT

OpenStudy (anonymous):

@rebeccaskell94 is it okay

OpenStudy (anonymous):

What is okay?

OpenStudy (anonymous):

my role up there